Researchers associated with South Miami Hospital (SMH) developed a new experimental laser treatment for heart patients. Its development team and the physicians who use the laser consider it to be a lifesaving advance. It should be noted that the physicians who are touting the laser hold a significant stake in the company that produces the laser. To offer a substitute for a balloon angioplasty to treat heart blockages, the experimental laser was developed at a cost of $250,000. SMH estimates that it will cost $20,000 to install the laser. The procedure requires a nurse at $50 per hour, a technician at $30 per hour, and a physician who is paid $750 per hour. Patients are billed $3,000 for the procedure compared to $1,500 for the traditional balloon treatment. Is it proper for physicians to recommend this procedure when they have a vested interest in its usage?
Answer #1

Total fixed cost = 250000+20000 = 270000

variable cost =50+30+750+1500(opportunity cost) = 2330/Hour

Income after new system= 3000/ hour

Contribution = Income per unit - variable cost

= 3000-2330

= 670 per hour

Break even point = Fixed cost / contribution per hour

= 270000/670

= 2481.48 hours

If doctor utilise this machine for 2481.48 hours he will be indifferent between the decision of install and not to install if he utilise machine for more than 2481.48 hours in that case they should go for New procedure.

I hope my efforts will be fruitful to you...?